Part A A weight lifter can bench press 0.29 kg. How many milligrams (mg) is this? ○ 2.9 × 107 mg 2.9 x 105 mg 29 108 mg 2.9 x 106 mg Submit
Answer #1

Bench press m = 0.29 kg 10° g 0- mg 1.0kg人1.0g m= 0.29×100 mg m= 2.9×105 mg